a jar contains 40 coins consisting of dimes and quarters and having a total value of $4.90. how many of each kind of coin are there?
Answer by checkley71(8403)

D+Q=40 & 10D+25Q=490 THUS D=40-Q OR 10(40-Q)+25Q=490 OR 400-10Q+25Q=490 OR
15Q=90 OR Q=90/15 OR Q=6 & D+6=40 OR D=34 25*6+34*10=490 OR 150+340=490 0R 490=490